you want to make a square-bottomed, open top box with a height of three inches and a volume of approximately 60 inches. you will take a square piece of metal and cut three-inch squares from each corner. you will then fold up the edges and weld them together to form the box. what should be the dimensions of the original piece of metal, to the nearest tenth of an inch?
Answer by mananth(16946)

you want to make a square-bottomed, open top box with a height of three inches and a volume of approximately 60 inches. you will take a square piece of metal and cut three-inch squares from each corner. you will then fold up the edges and weld them together to form the box. what should be the dimensions of the original piece of metal, to the nearest tenth of an inch?
...
let the length of side of the metal sheet be x inches.
you cut 3 in. square corners
so you reduce the length of the sheet by 6 inches.
Length of the side of the box = (x-6) inches.
Volume V = L*W*H( in this case L=W)
Volume = 60 cu.in
(x-6)(x-6)*3 = 60
(x-6)^2*3=60
/3
(x-6)^2= 20
take the square root
(x-6) = sqt 20
x-6 = 4.47
x= 6+4.47
x= 10.5 inches. The measure if the side of the sheet.
10.5 by 10.5 inches
..
m.ananth@hotmail.ca